OpenStudy (anonymous):

what is the solution of |2x-1|=5?

OpenStudy (anonymous):

Turn it into 2 separate equations and solve each one separately. 2x - 1 = 5 and 2x - 1 = -5

OpenStudy (jdoe0001):

\(\bf |2x-1|=5\implies \begin{cases} +(2x-1)=5\\ \quad \\ \bf -(2x-1)=5 \end{cases}\)

OpenStudy (anonymous):

Anytime you have an absolute value like that isolated on one side of the equation, you can turn it into 2 equations like that.. one is positive on the other side , and the other is negative on the other side.